Find an equation of the line having the given slope and containing the given point.
m=4/5, (4,-9)

Answer by Mathtut(3670)

y+9=(4/5)(x-4)
:
y+9=(4/5)x-16/5
:
5y+45=4x-16.........multiplied all terms by 5
:
4x-5y=61